# Interpreting Event-Studies from Recent Difference-in-Differences Methods

### v1

@@ Abstract
This note discusses the interpretation of event-study plots produced by recent difference-in-differences methods. I show that even when specialized to the case of non-staggered treatment timing, the default plots produced by software for three of the most popular recent methods (de Chaisemartin and D’Haultfœuille, 2020; Callaway and Sant’Anna, 2021; Borusyak et al., 2024) do not match those of traditional two-way fixed effects (TWFE) event-studies: the new methods may show a kink or jump at the time of treatment even when the TWFE event-study shows a straight line. This difference stems from the fact that the new methods construct the pre-treatment coefficients asymmetrically from the post-treatment coefficients. As a result, visual heuristics for analyzing TWFE event-study plots should not be immediately applied to those from these methods. I conclude with practical recommendations for constructing and interpreting event-study plots when using these methods.
@@ 1 Introduction
Modern difference-in-differences (DiD) analyses typically show an event-study plot that allows the researcher to evaluate differences in trends between the treated and comparison groups both before and after the treatment. [^1] Roth (2022) identifies 70 recent papers in top economics journals displaying such plots. Event-study plots allow the researcher to perform “visual inference” to determine whether the treated group’s trends appear to have deviated from the comparison group’s trends right around the time of treatment. An event-study plot is generally considered more convincing if there is a discontinuity or kink in the coefficients close to the treatment date, so that the pattern seen cannot simply be explained by a linear violations of parallel trends (Rambachan and Roth, 2023).
@@
Until recently, event-study plots were created by running a dynamic two-way fixed effects (TWFE) regression specification including indicators for time relative to treatment. A recent literature surveyed in de Chaisemartin and D’Haultfœuille (2022) and Roth et al. (2023) has illustrated that TWFE specifications may be difficult to interpret in settings with heterogeneous treatment effects, even under a parallel trends assumption. Most pertinently, Sun and Abraham (2021) showed that the event-study coefficient for relative time $r$ from a dynamic TWFE specification may be “contaminated” by treatment effects at a different relative time $r^{\prime}$. These issues with TWFE specifications have motivated the development of a variety of new DiD methods that yield more interpretable estimands in settings with staggered treatment timing and treatment effect heterogeneity. Many of the new methods allow for the construction of an event-study plot, and so it may be tempting to simply interpret the event-study produced by these new methods analogously to those from traditional dynamic TWFE specifications in the non-staggered setting.
@@
The main point of this note is that the default event-study plots produced by software for three of the most popular recent methods should not be interpreted in the same way as traditional dynamic TWFE specifications. Indeed, when specialized to the case with a common treatment date, the event-study plots from de Chaisemartin and D’Haultfœuille (2020, henceforth, dCDH), [^2] More precisely, the default plot produced by the R package for dCDH does not match the standard TWFE output; however, the default plot produced by the Stata package does. The authors have indicated to me that an updated version of the R package in forthcoming, which will match the output of the Stata package. Callaway and Sant’Anna (2021, henceforth, CS) and Borusyak et al. (2024, henceforth, BJS) do not match those from a dynamic TWFE specification. The plots from these new methods may show a kink or jump at the time of treatment even when there is no treatment effect and the violation of parallel trends is the same in all periods, so that the traditional dynamic TWFE specification shows a straight line. The typical heuristics for visual inference developed based on dynamic TWFE specifications will thus be misleading when applied to these new estimators.
@@
The remainder of this note is structured as follows. In Section 2, I provide a simple simulation illustrating how dynamic TWFE, dCDH, CS, and BJS produce very different event-study plots on the same data without staggered treatment timing. In Section 3, I provide a mathematical explanation showing that these differences arise because these estimators construct the pre-treatment and post-treatment event-study coefficients asymmetrically. Section 4 concludes with some recommendations.
@@ 2 A simulation comparing the event-studies in a non-staggered setting
I begin by illustrating the differences between the event-studies produced by TWFE and the newer methods with a simulation of an extremely simple non-staggered DiD setting. [^3] Code for the results shown in this section is available on Github, https://github.com/jonathandroth/HetEventStudies/tree/master. Units are either treated beginning at period $t=1$ or never-treated. We observe outcomes from period $t=-15$ to period $t=10$. There are no treatment effects in any period, so that $Y_{it}(1)=Y_{it}(0)=Y_{it}$. However, parallel trends is violated: the treated group’s outcome increases linearly relative to the control groups outcome in all periods. In particular, $Y_{it}=\gamma\cdot t+\epsilon_{it}$ for units in the treated group and $Y_{it}=\epsilon_{it}$ for units in the control group, where $\epsilon_{it}$ are $iid$ standard normal draws. Thus, the treated group’s outcome grows by $\gamma$ more than the control group’s outcome on average in each period. For the illustrative simulation, I set $\gamma=0.5$.
Figure 1 shows event-studies for a single simulation draw from this DGP when there are 100 total units and half of them are treated. Panel (a) shows the results of a traditional dynamic TWFE specification of the form
@@
where $D_{i}$ is an indicator for treatment and $\alpha_{i}$ and $\lambda_{t}$ are unit and period fixed effects. Panel (b) shows the results for the event-study generated by the CS method, using the default settings in the did package in R. Panel (c) shows the event-study plot generated using the dCDH method, using the default settings in the did_multiplegt R package. [^4] See FN 2 regarding differences between the Stata and R packages. Panel (d) shows the event-study plot generated by the BJS method, using the default settings in the did_imputation Stata package. [^5] All packages were installed between January 16-18, 2024.
Despite using identical data, the four methods produce very different event-study plots. The dynamic TWFE specification shows an approximately linear pre-trend that continues into the post-treatment period, with no apparent break in trend at the treatment date. By contrast, the CS and dCDH event-study plots have relatively flat pre-treatment coefficients and then exhibit a kink at the time of the treatment, sloping upwards afterwards. One can verify that in this example the coefficient estimates are numerically the same for CS and dCDH. [^6] The standard errors differ slightly owing to different implementations of the bootstrap. The BJS event-study appears to show an upward pre-trend, but then exhibits a sharp downward jump at the treatment date.
(a) Dynamic TWFE
(b) CS
(c) dCDH
(d) BJS
Figure 1: Comparison of event-study plots in a non-staggered setting
@@ 3 Mathematical Examination
In this section, I show that the differences between the event-plots created by dynamic TWFE and the newer estimators stems from the asymmetric way that the newer estimators construct the pre-treatment and post-treatment coefficients.
I consider a slightly generalized version of the set-up in the simulation section, in which we observe panel data for periods $t=-\underaccent{\bar}{T},...,\bar{T}$. Treated units denoted by $D_{i}=1$ begin treatment at period 1, whereas comparison units ($D_{i}=0$) are untreated in all periods. The observed outcome is $Y_{it}=D_{i}Y_{it}(1)+(1-D_{i})Y_{it}(0)$. For some of our results, we will consider the setting in the simulation where $Y_{it}(1)=Y_{it}(0)=\gamma\cdot t+\epsilon_{it}$ where $E[\epsilon_{it}\mid D_{i}]=0$.
@@ 3.1 Dynamic TWFE
It is straightforward to show that the event-study coefficients are given by
@@
where $\hat{E}$ denotes the sample average. That is, $\hat{\beta}_{r}$ shows a 2-group, 2-period DiD estimate comparing the $D_{i}=1$ and $D_{i}=0$ groups between periods $r+1$ and 00. Note that the formula for $\hat{\beta}_{r}$ is symmetric between pre-treatment and post-treatment periods—i.e. the same regardless of whether $r$ is greater than or less than zero.
Under mild regularity conditions, the sample means converge to population means, and thus $\hat{\beta}_{r}$ is consistent for
@@
3.2 CS and dCDH
I now turn to the construction of the CS and dCDH event-study estimates, which recall are numerically equivalent in this simple non-staggered setting. [^7] Specifically, I am referring to the dCDH estimates in R; see FN 2. CS/dCDH construct event-studies by aggregating DiD comparisons of treated and not-yet-treated units. However, by default, the construction is asymmetric across the pre-treatment and post-treatment periods. In particular, the pre-treatment coefficients average “short-differences”, i.e. comparisons of consecutive periods, whereas the post-treatment coefficients are “long-differences”, i.e. comparisons relative to the period before treatment.
@@ 3.2 CS and dCDH
$\hat{\beta}_{r}^{CS/dCDH}=\begin{cases}\left(\hat{E}[Y_{i,r+1}\mid D_{i}=1]-\hat{E}[Y_{i,r+1}\mid D_{i}=0]\right)-\left(\hat{E}[Y_{i,r}\mid D_{i}=1]-\hat{E}[Y_{i,r}\mid D_{i}=0]\right)&\text{ if }r<0\\ \left(\hat{E}[Y_{i,r+1}\mid D_{i}=1]-\hat{E}[Y_{i,r+1}\mid D_{i}=0]\right)-\left(\hat{E}[Y_{i0}\mid D_{i}=1]-\hat{E}[Y_{i0}\mid D_{i}=0]\right)&\text{ if}r\geq 0\end{cases}$
Note that for $r<0$, the comparison period is period $r-1$, whereas for $r\geq 0$ the comparison period is always period 00. Under mild regularity conditions, it follows that $\hat{\beta}_{r}^{CS/dCDH}$ is consistent for
$\beta_{r}^{CS/dCDH}=\begin{cases}\left(E[Y_{i,r+1}\mid D_{i}=1]-E[Y_{i,r+1}\mid D_{i}=0]\right)-\left(E[Y_{i,r}\mid D_{i}=1]-E[Y_{i,r}\mid D_{i}=0]\right)&\text{ if }r<0\\ \left(E[Y_{i,r+1}\mid D_{i}=1]-E[Y_{i,r+1}\mid D_{i}=0]\right)-\left(E[Y_{i0}\mid D_{i}=1]-E[Y_{i0}\mid D_{i}=0]\right)&\text{ if }r\geq 0\end{cases}$
@@
$\beta_{r}^{CS/dCDH}=\begin{cases}\gamma&\text{ if }r<0\\ \gamma\cdot(r+1)&\text{ if }r\geq 0\end{cases}$
so the population version of the CS/dCDH event-study coefficients show a kink at zero if $\gamma\neq 0$.
@@ 3.3 BJS
Like the CS/dCDH event-study approach, the BJS event-study is also constructed in a way that is asymmetric between the pre-treatment and post-treatment periods. For the post-treatment effects, they use what they call an imputation approach: they fit a TWFE model using only untreated $(i,t)$ pairs, then form individual treatment effect estimates of the form $\hat{\tau}_{it}=Y_{it}-\hat{Y}_{it}$, where $\hat{Y}_{it}$ is the prediction from the TWFE model. They then average the $\hat{\tau}_{it}$ for units at a given lag from treatment. For the pre-treatment model, they simply run a dynamic TWFE specification using only $(i,t)$ pairs that are untreated, with dynamic indicators for the number of periods until treatment (with the earliest pre-treatment period normalized to zero). [^8] The Stata package requires the researcher to specify how many pre-treatment coefficients they would like to calculate, and all earlier periods are used as the omitted category. The omitted category is thus the earliest period if one requests $\underaccent{\bar}{T}$ pre-treatment coefficients.
In our special case of non-staggered treatment timing, the post-treatment event-study coefficients for BJS correspond to simple DiDs between period $r+1$ and the average outcome in the pre-treatment period. By contrast, the pre-treatment coefficients correspond to DiDs between period $r+1$ and the earliest period, period $-\underaccent{\bar}{T}$. Specifically, let $\bar{Y}_{i}^{pre}=\frac{1}{\underaccent{\bar}{T}+1}\sum_{t=-\underaccent{\bar}{T}}^{0}Y_{it}$ be $i$’s average outcome in periods $t\leq 0$. Then we have
$\hat{\beta}_{r}^{BJS}=\begin{cases}\left(\hat{E}[Y_{i,r+1}\mid D_{i}=1]-\hat{E}[Y_{i,r+1}\mid D_{i}=0]\right)-\left(\hat{E}[Y_{i,-\underaccent{\bar}{T}}\mid D_{i}=1]-\hat{E}[Y_{i,-\underaccent{\bar}{T}}\mid D_{i}=0]\right)&\text{ if }r<0\\ \left(\hat{E}[Y_{i,r+1}\mid D_{i}=1]-\hat{E}[Y_{i,r+1}\mid D_{i}=0]\right)-\left(\hat{E}[\bar{Y}_{i}^{pre}\mid D_{i}=1]-\hat{E}[\bar{Y}_{i}^{pre}\mid D_{i}=0]\right)&\text{ if }r\geq 0\end{cases}$
This makes clear that the construction is asymmetric, with $\hat{\beta}_{r}^{BJS}$ using a comparison to period $-\underaccent{\bar}{T}$ for the pre-treatment coefficients and to the average of the pre-period for the post-treatment coeffiients.
@@
$\beta_{r}^{BJS}=\begin{cases}\left(E[Y_{i,r+1}\mid D_{i}=1]-E[Y_{i,r+1}\mid D_{i}=0]\right)-\left(E[Y_{i,-\underaccent{\bar}{T}}\mid D_{i}=1]-E[Y_{i,-\underaccent{\bar}{T}}\mid D_{i}=0]\right)&\text{ if }r<0\\ \left(E[Y_{i,r+1}\mid D_{i}=1]-E[Y_{i,r+1}\mid D_{i}=0]\right)-\left(E[\bar{Y}_{i}^{pre}\mid D_{i}=1]-E[\bar{Y}_{i}^{pre}\mid D_{i}=0]\right)&\text{ if }r\geq 0\end{cases}$

### v2

@@ Abstract
This note discusses the interpretation of event-study plots produced by recent difference-in-differences methods. I show that even when specialized to the case of non-staggered treatment timing, the default plots produced by software for several of the most popular recent methods do not match those of traditional two-way fixed effects (TWFE) event-studies. The plots produced by the new methods may show a kink or jump at the time of treatment even when the TWFE event-study shows a straight line. This difference stems from the fact that the new methods construct the pre-treatment coefficients asymmetrically from the post-treatment coefficients. As a result, visual heuristics for evaluating violations of parallel trends using TWFE event-study plots should not be immediately applied to those from these methods. I conclude with practical recommendations for constructing and interpreting event-study plots when using these methods.
@@ 1 Introduction
Modern difference-in-differences (DiD) analyses typically report an event-study plot that shows the differences in trends in the outcome between the treated and comparison groups both before and after the treatment date.111roth_pre-test_2021 identifies 70 recent papers in top economics journals displaying such plots. Event-study plots serve two purposes. First, if the DiD design is valid, they allow the reader to evaluate the dynamic patterns of treatment effects following the implementation of the policy of interest. Second, event-study plots provide evidence about the validity of the research design. They enable the reader to perform “visual inference” to determine to what extent the estimated treatment effects may be driven by violations of the parallel trends assumption. An event-study plot is generally considered more convincing evidence in favor of a non-zero treatment effect if there is a discontinuity or kink in the coefficients close to the treatment date, so that the pattern seen cannot simply be explained by a smooth violation of parallel trends (freyaldenhoven_visualization_2021; rambachan_more_2023).
@@
Until recently, event-study plots were typically created by estimating a dynamic two-way fixed effects (TWFE) regression specification, with fixed effects for unit and time and indicators for time relative to treatment (with the coefficient for the period directly before treatment typically normalized to zero). A recent literature surveyed in de_chaisemartin_two-way-survey_2021 and roth_whats_2023 has illustrated that such TWFE specifications may be difficult to interpret in settings with heterogeneous treatment effects and staggered treatment timing, even if a parallel trends assumption holds. Most pertinently, sun_estimating_2020 showed that the event-study coefficient for relative time $r$ from a dynamic TWFE specification may be “contaminated” by treatment effects at a different relative time $r^{\prime}$. This undermines the first purpose of event-study plots described above, since the post-treatment event-study coefficients may not correspond directly to the dynamics of the treatment effects. These issues with TWFE specifications have motivated the development of a variety of new DiD methods that allow for a more straightforward evaluation of dynamic treatment effects, assuming parallel trends holds, in settings with staggered timing and heterogeneous treatment effects. These new methods typically produce event-studies that also show pre-treatment coefficients. It would then seem natural to also use these plots for the second purpose of evaluating the extent to which the results could be explained by violations of parallel trends, using the same heuristics as with traditional dynamic TWFE specifications in the non-staggered setting.
@@
The main point of this note is that for the purposes of assessing violations of parallel trends, the default event-study plots produced by software for several of the most popular recent methods should not be interpreted in the same way as traditional dynamic TWFE specifications. Indeed, I show that these event-study plots do not match those of traditional dynamic TWFE specifications even when specialized to the setting with non-staggered treatment timing. This can be seen in Figure 1 below, which shows the event-studies produced by different methods when applied to the same simulated dataset in which there is non-staggered treatment timing. In the simulation (the details of which are described below) there is no effect of the treatment but parallel trends is violated: the outcomes for the treated group are on an upwards linear trend relative to the comparison group. As expected, the event-study plot produced using dynamic TWFE yields coefficients that lie close to a straight line. By contrast, the event-study plot produced by the R package for implementing the method of callaway_difference--differences_2020 shows a clear kink at the treatment date. [^2] As I describe below, the identical pattern also arises using a previous version of the R package for implementing the method of de_chaisemartin_two-way_2020. Subsequent to the release of the initial draft of this paper in January 2024, the R package for dCDH was updated to produce results matching the TWFE event-study. Meanwhile, the event-study plot produced by the Stata package for borusyak_revisiting_2024 shows a clear discontinuity at the treatment date. Given these starkly different results when applied to the same data-set with non-staggered treatment timing, it is clear that the typical heuristics for visual inference developed based on dynamic TWFE specifications will be misleading when applied to these new estimators.
@@
The remainder of this note is structured as follows. In Section 2, I provide a simple simulation illustrating how recent methods can produce very different event-study plots on the same data without staggered treatment timing. In Section 3, I provide a mathematical explanation showing that these differences arise because these estimators construct the pre-treatment and post-treatment event-study coefficients asymmetrically. Section 4 concludes with some recommendations and further discussion.
(a) Dynamic TWFE
(b) CS (R package)
(c) BJS (Stata package)
Figure 1: Comparison of event-study plots in a non-staggered setting
@@ 2 A simulation comparing the event-studies in a non-staggered setting
I begin by illustrating the differences between the event-studies produced by TWFE and some newer methods with a simulation of an extremely simple non-staggered DiD setting. [^3] Code for the results shown in this section is available on Github, https://github.com/jonathandroth/HetEventStudies/tree/master. Units are either treated beginning at period $t=1$ or never-treated. We observe outcomes from period t=15t=-15 to period $t=10$. There are no treatment effects in any period, so that $Y_{it}(1)=Y_{it}(0)=Y_{it}$. However, parallel trends is violated: the treated group’s outcome increases linearly relative to the control group’s outcome in all periods. In particular, $Y_{it}(0)=\gamma\cdot t+\epsilon_{it}$ for units in the treated group and $Y_{it}(0)=\epsilon_{it}$ for units in the control group, where $\epsilon_{it}$ are iidiid standard normal draws. Thus, the treated group’s outcome grows by $\gamma$ more than the control group’s outcome on average in each period. For the illustrative simulation, I set $\gamma=0.5$.
Figure 1 shows event-studies for a single simulation draw from this DGP when there are 100 total units and half of them are treated. Panel (a) shows the results of a traditional dynamic TWFE ordinary least-squares specification of the form
@@
where $D_{i}$ is an indicator for treatment and $\alpha_{i}$ and $\lambda_{t}$ are unit and period fixed effects. Panel (b) shows the results for the event-study generated by the CS method, using the default settings in the did package in R. Panel (c) shows the event-study plot generated by the BJS method, using the default settings in the did_imputation Stata package. [^4] All packages were re-installed using the latest versions as of October 31, 2025. I obtained similar results as of the initial draft of this paper in January 2024, with one exception owing to a software update described below (see Figure 2 and surrounding discussion).
As described in the introduction, despite using identical data, the three methods produce very different event-study plots. The dynamic TWFE specification shows an approximately linear pre-trend that continues into the post-treatment period, with no apparent break in trend at the treatment date. By contrast, the CS event-study plot has relatively flat pre-treatment coefficients and then exhibits a kink at the time of the treatment, sloping upwards afterwards. The BJS event-study appears to show an upward pre-trend, but then exhibits a sharp downward jump at the treatment date.
(a) dCDH (Previous R implementation)
(b) dCDH (Current R implementation)
(c) Wooldridge (etwfe package)
(d) SA (fixest implementation)
Figure 2: Comparison of event-studies from other methods in the non-staggered setting
Other estimators.
Although I focus on CS and BJS, two of the most popular recent methods, I now comment briefly on the event-study plots produced by other recent DiD methods. At the time of the release of the initial draft of this paper in January 2024, the DIDmultiplegt R package for implementing the method of de_chaisemartin_two-way_2020 produced identical point estimates to the CS event-study described above; see Figure 2. However, the most recent version of the package produces results that match the original TWFE event-study plot in my simple example. [^5] The dCDH package only reports as many pre-trends coefficients as there are post-treatment periods. Thus, it matches the TWFE event-study for the 10 latest pre-treatment periods but does not show the earlier pre-trends coefficients. Two other popular approaches are those of sun_estimating_2020 and wooldridge_two-way_2025. Both of these papers propose richer TWFE specifications to capture treatment effect heterogeneity in staggered settings. However, these TWFE specifications reduce to the usual dynamic TWFE event-study in (2) when specialized to the case with non-staggered treatment timing, and thus match panel (a) of Figure 1 by construction. This can be seen in panels (c) and (d) of Figure 2, which shows implementations of SA and W from the etwfe and fixest packages in R. [^6] For etwfe, we must specify the option cgroup = ‘‘never’’ in order to get pre-trends estimates. Two other popular approaches are those of liu_practical_2024 and gardner_two-stage_2021. Like BJS, these papers rely on an “imputation” approach for inferring counterfactual outcomes, although they construct the pre-treatment coefficients in the event-study differently. In Appendix Section A, I discuss the results produced by the packages for these papers (fect and did2s). Both of these methods also yield an event-study that shows a jump at the date of treatment.
@@ 3 Mathematical Examination
In this section, I show that the differences between the event-plots created by dynamic TWFE and CS and BJS stems from the asymmetric way that the newer estimators construct the pre-treatment and post-treatment coefficients.
I consider a slightly generalized version of the set-up in the simulation section, in which we observe panel data for periods $t=-\underaccent{\bar}{T},...,\bar{T}$. Treated units denoted by $D_{i}=1$ begin treatment at period 1, whereas comparison units ($D_{i}=0$) are untreated in all periods. The observed outcome is $Y_{it}=D_{i}Y_{it}(1)+(1-D_{i})Y_{it}(0)$. For some of our results, we will consider the setting in the simulation where $Y_{it}(1)=Y_{it}(0)=\gamma\cdot t\cdot D_{i}+\epsilon_{it}$ and $E[\epsilon_{it}\mid D_{i}]=0$.
@@ 3.1 Dynamic TWFE
It is straightforward to show that the event-study coefficients estimated by OLS are given by
@@
where $\hat{E}$ denotes the sample average. That is, $\hat{\beta}_{r}$ shows a 2-group, 2-period DiD estimate comparing the $D_{i}=1$ and $D_{i}=0$ groups between periods $r+1$ and 0. Note that the formula for $\hat{\beta}_{r}$ is symmetric between pre-treatment and post-treatment periods—i.e. the same regardless of whether $r$ is greater than or less than zero.
Under mild regularity conditions, the sample means converge in probability to population means, and thus $\hat{\beta}_{r}$ is consistent for
@@
3.2 CS
I now turn to the construction of the CS event-study estimates. CS constructs event-studies by aggregating DiD comparisons of treated and not-yet-treated units. However, by default, the construction is asymmetric across the pre-treatment and post-treatment periods. In particular, the pre-treatment coefficients are “short-differences”, i.e. comparisons of consecutive periods, whereas the post-treatment coefficients are “long-differences”, i.e. comparisons relative to the period before treatment.
@@
$\hat{\beta}_{r}^{CS}=\begin{cases}\left(\hat{E}[Y_{i,r+1}\mid D_{i}=1]-\hat{E}[Y_{i,r+1}\mid D_{i}=0]\right)-\left(\hat{E}[Y_{i,r}\mid D_{i}=1]-\hat{E}[Y_{i,r}\mid D_{i}=0]\right)&\text{ if }r<0\\ \left(\hat{E}[Y_{i,r+1}\mid D_{i}=1]-\hat{E}[Y_{i,r+1}\mid D_{i}=0]\right)-\left(\hat{E}[Y_{i0}\mid D_{i}=1]-\hat{E}[Y_{i0}\mid D_{i}=0]\right)&\text{ if }r\geq 0.\end{cases}$
Note that for pre-treatment coefficients ($r<0$), we compare outcomes in consecutive periods (periods $r+1$ vs $r$), whereas for post-treatment coefficients ($r\geq 0$), we compare outcomes over longer horizons (periods $r+1$ vs 0). Under mild regularity conditions, it follows that $\hat{\beta}_{r}^{CS}$ is consistent for
$\beta_{r}^{CS}=\begin{cases}\left(E[Y_{i,r+1}\mid D_{i}=1]-E[Y_{i,r+1}\mid D_{i}=0]\right)-\left(E[Y_{i,r}\mid D_{i}=1]-E[Y_{i,r}\mid D_{i}=0]\right)&\text{ if }r<0\\ \left(E[Y_{i,r+1}\mid D_{i}=1]-E[Y_{i,r+1}\mid D_{i}=0]\right)-\left(E[Y_{i0}\mid D_{i}=1]-E[Y_{i0}\mid D_{i}=0]\right)&\text{ if }r\geq 0.\end{cases}$
@@
$\beta_{r}^{CS}=\begin{cases}\gamma&\text{ if }r<0\\ \gamma\cdot(r+1)&\text{ if }r\geq 0\end{cases}$
so the population version of the CS event-study coefficients show a kink at zero if $\gamma\neq 0$.
@@ 3.3 BJS
Like the CS event-study approach, the BJS event-study is also constructed in a way that is asymmetric between the pre-treatment and post-treatment periods. For the post-treatment effects, they use what they call an imputation approach: they fit a TWFE regression using only untreated $(i,t)$ pairs, then form individual treatment effect estimates of the form $\hat{\tau}_{it}=Y_{it}-\hat{Y}_{it}$, where $\hat{Y}_{it}$ is the prediction from the TWFE regression. They then average the $\hat{\tau}_{it}$ for treated units at a given lag from treatment. For the pre-treatment coefficients, they simply run a dynamic TWFE regression using only $(i,t)$ pairs that are untreated, with dynamic indicators for the number of periods until treatment (with the earliest pre-treatment period normalized to zero). [^7] The Stata package requires the researcher to specify how many pre-treatment coefficients they would like to calculate, and all earlier periods are used as the omitted category. The omitted category is thus the earliest period if one requests $\underaccent{\bar}{T}$ pre-treatment coefficients.
In our special case of non-staggered treatment timing, the post-treatment event-study coefficients for BJS correspond to simple DiDs between period $r+1$ and the average outcome in the pre-treatment period. By contrast, the pre-treatment coefficients correspond to DiDs between period $r+1$ and the earliest period, $\underaccent{\bar}{T}$. Specifically, let $\bar{Y}_{i}^{pre}=\frac{1}{\underaccent{\bar}{T}+1}\sum_{t=-\underaccent{\bar}{T}}^{0}Y_{it}$ be $i$’s average outcome in periods $t\leq 0$. Then we have
$\hat{\beta}_{r}^{BJS}=\begin{cases}\left(\hat{E}[Y_{i,r+1}\mid D_{i}=1]-\hat{E}[Y_{i,r+1}\mid D_{i}=0]\right)-\left(\hat{E}[Y_{i,-\underaccent{\bar}{T}}\mid D_{i}=1]-\hat{E}[Y_{i,-\underaccent{\bar}{T}}\mid D_{i}=0]\right)&\text{ if }r<0\\ \left(\hat{E}[Y_{i,r+1}\mid D_{i}=1]-\hat{E}[Y_{i,r+1}\mid D_{i}=0]\right)-\left(\hat{E}[\bar{Y}_{i}^{pre}\mid D_{i}=1]-\hat{E}[\bar{Y}_{i}^{pre}\mid D_{i}=0]\right)&\text{ if }r\geq 0.\end{cases}$
This makes clear that the construction is asymmetric, with $\hat{\beta}_{r}^{BJS}$ using a comparison to period $\underaccent{\bar}{T}$ for the pre-treatment coefficients and to the average of the pre-treatment periods for the post-treatment coefficients.
@@
$\beta_{r}^{BJS}=\begin{cases}\left(E[Y_{i,r+1}\mid D_{i}=1]-E[Y_{i,r+1}\mid D_{i}=0]\right)-\left(E[Y_{i,-\underaccent{\bar}{T}}\mid D_{i}=1]-E[Y_{i,-\underaccent{\bar}{T}}\mid D_{i}=0]\right)&\text{ if }r<0\\ \left(E[Y_{i,r+1}\mid D_{i}=1]-E[Y_{i,r+1}\mid D_{i}=0]\right)-\left(E[\bar{Y}_{i}^{pre}\mid D_{i}=1]-E[\bar{Y}_{i}^{pre}\mid D_{i}=0]\right)&\text{ if }r\geq 0.\end{cases}$